OpenStudy (anonymous):

27x^2-9x^3-18x^4 ------------------ divide thank you!! -18x^3

OpenStudy (mertsj):

\[\frac{27x^2-9x^3-18x^4}{-18x^3}=\frac{27x^2}{-18x^3}-\frac{9x^3}{-18x^3}-\frac{18x^4}{-18x^3}\]

OpenStudy (mertsj):

Simplify each fraction.

OpenStudy (anonymous):

so is it -3/2x^-1 -1/2x -x and thank you!!

OpenStudy (mertsj):

Should be -3/2x^-1+1/2+x

OpenStudy (anonymous):

thank you so much!!